OpenStudy (anonymous):

Out of 1,100 discs tested, 13 are defective. Estimate the number of defective discs in a batch of 41,000. A. about 485 discs B. about 970 discs C. about 4,850 discs D. about 49 discs

OpenStudy (anonymous):

13/1100 = x / 41000 x = 41000 x 13 /1100 answer is a